a brass ring has a diameter of 20cm when placed in melting ice at 0°Celsius. what will it's diameter be if it is placed in a boiling water?
1
Expert's answer
2020-03-02T10:12:55-0500
"d=d_0(1+\alpha (T-T_0))"

"d=20(1+19\cdot 10^{-6} (100-0))=20.038\ cm"